OpenStudy (anonymous):

I need help, but i also want to learn how, not just answers. :) its graphing quadratic equations

OpenStudy (asnaseer):

hi - what specifically do you need help on?

OpenStudy (anonymous):

okay let me right down the equation :) just a sec...

OpenStudy (anonymous):

x^2=-1

OpenStudy (anonymous):

how would i go about graphing this?

OpenStudy (asnaseer):

ok, so first lets rearrange this to get:\[x^2+1=0\]

OpenStudy (anonymous):

okay, so you have to make it equal to zero to graph?

OpenStudy (anonymous):

oh! sorry, its -10

OpenStudy (anonymous):

my bad

OpenStudy (asnaseer):

now the graph of the function:\[y=x^2+1\] will show you how the function \(x^2+1\)changes, and, when it equals zero, this will give you a solution to your equation \(x^2+1=0\)

OpenStudy (asnaseer):

thats ok, we just change it to \(y=x^2+10\)

OpenStudy (anonymous):

okay :)

OpenStudy (asnaseer):

do you know how to graph this function?

OpenStudy (anonymous):

no, I'm not sure

OpenStudy (anonymous):

its says sketch graph and axis of symmetry

OpenStudy (asnaseer):

ok, let me try and draw it for you and then explain how I got there

OpenStudy (anonymous):

For such problems, I would like to typically create a table and then plot those values on a graph. Like the following: